OpenStudy (anonymous):

Please Help... A robot travels at a speed of 1.35 m/sec at a direction of 45.0 degrees for 305 seconds. It then travels at a speed of 1.50 m/sec at a direction of 140.0 degrees for 852 seconds. What is the total displacement at the end of the second maneuver? First, give the y component of the first displacement vector calculated in the problem.

OpenStudy (anonymous):

1308.08m u can use the law of cosines c^2=a^2+b^2-2ab cos (c) where cos(c)=cos(85degree)

OpenStudy (anonymous):

Is 1308.08 the y component or the total displacement?

OpenStudy (anonymous):

the total

OpenStudy (anonymous):

Firstly, I'm assuming that the degrees are calculated from a absolute degree frame. i.e. North is 000, East is 090. If not, please tell. The first , y-component, can be calc'ed by this: \(S_y=u_y t + \frac{1}{2} a_y t^2\) Now, since \(u_y\) will be \(u sin 45^o\)

OpenStudy (anonymous):

Then, a_y= 0, t=305, sub it all in to find the \(S_y\) which is the y displacement.

OpenStudy (anonymous):

Do you need further help?